OpenStudy (anonymous):

Find the flux of the vector field H = 4xy i  + z^2 j  + 12y k  out of the closed box 0 ≤ x ≤ 3, 0 ≤ y ≤ 2, 0 ≤ z ≤ 8.

ganeshie8 (ganeshie8):

use divergence thm

OpenStudy (dan815):

hi baby

OpenStudy (anonymous):

I got div H= y so flux is \[\int\limits_{0}^{8}\int\limits_{0}^{2}\int\limits_{0}^{3}ydxdydz\]

OpenStudy (anonymous):

equals 48 right but its making it wrong

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

what am I doing wrong?

ganeshie8 (ganeshie8):

you're missing a 4 i think..

OpenStudy (dan815):

yeah!

OpenStudy (dan815):

120+32=152

OpenStudy (anonymous):

152?

OpenStudy (dan815):

oops

OpenStudy (dan815):

160+32=192

OpenStudy (anonymous):

48 * 4=192

OpenStudy (dan815):

:)

OpenStudy (anonymous):

so the div H=4y right?

OpenStudy (anonymous):

thanks guys

ganeshie8 (ganeshie8):

divergence(Pi + Qj + Rk) = P_x + Q_y + R_z

ganeshie8 (ganeshie8):

divergence(4xy i  + z^2 j  + 12y k ) = 4y + 0 + 0 = 4y

ganeshie8 (ganeshie8):

refresh to see that mess

OpenStudy (anonymous):

I see it now